OpenStudy (anonymous):

Find the radius of a circle that has circumference 12 more than its diameter.

OpenStudy (nottim):

What's the formula for circumference?

OpenStudy (anonymous):

2\[\pi\]r

OpenStudy (anonymous):

d=2(pi)r

OpenStudy (anonymous):

i mean c=2(pi)r

OpenStudy (anonymous):

\[d + 12 = d \pi\] since d = 2r \[(d + 12)\div d = \pi\] \[1 + (12 \div d) = \pi\] \[12 \div d = \pi - 1\] \[1 \div d = (\pi - 1) \div 12\] invert \[d = 12 \div (\pi - 1)\] \[d = 12 \div2.14\] d = 5.6 C = 5.6 * 3.14 = 17.6 which is 12 more than the diameter of 5.6 (rounded)